A trench is 180 feet long and 12 feet deep,7 feet wide at the top and 4 feet at the bottom.How many cubic yards of earth have been removed?
Answer by checkley77(12844)

180*12*(7+4)/2=2,160*11/2=11,880 cubic feet of earth.
11,880/27=40 cubic yards of earth.